OpenStudy (anonymous):

If cos s = 1/2 and 3pi/2 < s < 2pi, Find Cosecant S.

OpenStudy (wwe123):

270<s<360 cos s = 330 cos s = \[\frac{ 11\pi }{ 6 }\]

OpenStudy (anonymous):

my professor put the answer for cosecant = -\[-\frac{ 2 }{ \sqrt{3} }\]

OpenStudy (anonymous):

negative

OpenStudy (wwe123):

oh

OpenStudy (wwe123):

cosecant = 330 cosecant= -30 cosine = 1/cosecant cosine = 30 = \[\frac{ \sqrt{3} }{2 }\] cosine = -\[-\frac{ \sqrt{3} }{2}\] cosecant=\[-\frac{ 2 }{\sqrt{3} }\]

OpenStudy (anonymous):

csc is not 1/cosine!!!!! csc x = 1/sin x

OpenStudy (wwe123):

ok

OpenStudy (wwe123):

oh sorry

OpenStudy (wwe123):

just find cosec 330

OpenStudy (anonymous):

Google the unit circle and you need to find csc of 300 or csc 5pi/3

OpenStudy (anonymous):

csc is hypotenuse over opposite side

OpenStudy (anonymous):

cos s = 1/2 is when s = 300

OpenStudy (wwe123):

yes @ChrisS

OpenStudy (anonymous):

cos is the x interval

OpenStudy (anonymous):

alright i get it